OpenStudy (anonymous):

Solve the system by graphing. -5x - y = -11 2x - y = 10

OpenStudy (anonymous):

do you have the answer choices?

OpenStudy (anonymous):

here they are

OpenStudy (anonymous):

@mark_o.

OpenStudy (anonymous):

-5x - y = -11 ------> mult by -1 to get 5x+y+11 2x - y = 10 2x - y = 10 add --------------- 7x +0 =21 is x= 3 ? if it is solve now for y 2x - y = 10 sub x=3 2(3)-y=10 is y= -4 ?.... if yes is the answer would be like (x,y)=(3,-4) ?........

OpenStudy (anonymous):

@ algebra2sucks ...:D

OpenStudy (anonymous):

thank yooou :)